OpenStudy (anonymous):

Point E is located at (2, −3), and point F is located at (−2, −1). Find the y value for the point that is 3 over 4 the distance from point E to point F. 1.75 0.5 −2 −1.5

OpenStudy (aum):

The y value of the point 3/4th the distance of EF from E is: \(\large y_1 + \frac 34 * (y_2-y_1) = -3 + \frac 34 * (-1-(-3)) = ?\)

OpenStudy (anonymous):

-1.5

OpenStudy (aum):

correct.